# Investigating Context-Faithfulness in Large Language Models: The Roles of Memory Strength and Evidence Style

### v1

@@ Abstract
Retrieval-augmented generation (RAG) improves Large Language Models (LLMs) by incorporating external information into the response generation process. However, how context-faithful LLMs are and what factors influence LLMs’ context-faithfulness remain largely unexplored. In this study, we investigate the impact of memory strength and evidence presentation on LLMs’ receptiveness to external evidence. We introduce a method to quantify the memory strength of LLMs by measuring the divergence in LLMs’ responses to different paraphrases of the same question, which is not considered by previous works. We also generate evidence in various styles to evaluate the effects of evidence in different styles. Two datasets are used for evaluation: Natural Questions (NQ) with popular questions and popQA featuring long-tail questions. Our results show that for questions with high memory strength, LLMs are more likely to rely on internal memory, particularly for larger LLMs such as GPT-4. On the other hand, presenting paraphrased evidence significantly increases LLMs’ receptiveness compared to simple repetition or adding details.
Investigating Context-Faithfulness in Large Language Models: The Roles of Memory Strength and Evidence Style
Yuepei Li,   Kang Zhou,   Qiao Qiao,   Bach Nguyen,   Qing Wang,   Qi Li Department of Computer Science, Iowa State University, Ames, Iowa, USA {liyp0095, kangzhou, qqiao1, ntbach, qingwang, qli}@iastate.edu
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Fan et al. (2024); Zhao et al. (2023) has gained increasing popularity as it improves the performance of Large Language Models (LLMs) by integrating external information during the generation process, particularly when the model’s internal knowledge is insufficient or outdated Bianchini et al. (2024); Procko (2024); Siriwardhana et al. (2023); Vakayil et al. (2024); Wang et al. (2024); Jeong (2023). It raises the importance of the study of how context-faithful LLMs are, as RAG relies on effectively incorporating retrieved external information into the generated responses to ensure accuracy and contextual relevance. In this study, we explore whether LLMs are context-faithful when encountering external information, particularly when that information conflicts with the LLMs’ internal memory.
@@
To investigate the issue of context-faithfulness, one approach Longpre et al. (2021); Chen et al. (2022) is to create conflicts through entity substitution. Longpre et al. (2021) concluded that models frequently rely on their parametric knowledge, generating responses not found in the provided evidence. In contrast, Chen et al. (2022) found that LLMs rely almost exclusively on evidence passages rather than internal memory when multiple evidence passages are presented. Another approach Xie et al. (2024); Jin et al. (2024) involves generating counter-memory evidence with LLMs, and these studies have shown that LLMs are generally receptive to external evidence as long as it is coherent.
These methods, however, overlook some important aspects of the task. On one hand, LLMs have different memory strengths for different knowledge, and their behavior towards external evidence may vary significantly depending on how strongly it resistant to that knowledge. On the other hand, the characteristics of the evidence, such as evidence length, expressions, and detailed information involved, can influence how persuasive the evidence is. The behavior of LLMs is also expected to vary with evidence of different persuasiveness. Overlooking these factors could result in conflicting conclusions.
@@
To address these issues, we introduce a method to quantify the memory strength of LLMs. We also generate evidence in various styles to evaluate the different effects of evidence in different styles. Inspired by Zhao et al. (2024), we assess memory strength by measuring the divergence in LLMs’ responses to different paraphrases of the same question. Intuitively, an LLM demonstrates high memory strength when it consistently provides the same answer across all paraphrased versions of a question. For evidence style, We classify it into direct and indirect forms: direct evidence provides a straightforward answer to the question, while indirect evidence incorporates additional details to support the answer. Through these methods, we analyze the relationship between context-faithfulness and LLM memory strength, and we explore the impact of different evidence styles on context-faithfulness. Our conclusions are as follows:
@@
  • The receptiveness of LLMs to external evidence is strongly correlated with memory strength to the question. We observed this relationship both across different datasets and different LLMs. Contrary to the findings of Xie et al. (2024); Jin et al. (2024) that LLMs are highly receptive (less than 5%) to external evidence when it is coherent, we find that the probability of the model relying on its parametric memory is non-negligible for questions that the LLMs have a strong memory. For example, GPT-4, which has strong memory on the NQ dataset, reaches almost 50% memory answer ratio when answering questions from the NQ dataset.
@@
  • The style of the evidence plays an important role in determining the LLM’s receptiveness to external information. Our research demonstrates that presenting the LLM with multiple paraphrases of the same evidence substantially increases its receptiveness. This approach outperforms simple repetition of the evidence and is more effective than adding additional details to the evidence. These findings provide valuable insights for work in research of RAG and context faithful, highlighting the importance of prioritizing LLM receptiveness in both model training and application.
@@ 2.1 Context Faithfulness of LLM
To update static factual knowledge Lazaridou et al. (2021); Karpukhin et al. (2020); Kasai et al. (2023) in LLMs, the retrieval-based method has been introduced to involve external information to LLMs Lazaridou et al. (2022); Izacard et al. (2024); Khattab et al. (2022); Santhanam et al. (2022); Gao and Callan (2022). However, these methods can introduce knowledge conflicts between the introduced context and pre-existing memories from LLMs. LLMs often persist in relying on their pre-existing memories, overlooking newly provided contextual evidence Longpre et al. (2021). To address this, recent studies Neeman et al. (2023); Li et al. (2023) fine-tune LLMs on counterfactual contexts, where the original facts are replaced with counterfactual ones. Another work Zhou et al. (2023) proposes a novel approach using prompting to improve context faithfulness in LLMs without additional fine-tuning.
A related area of research focuses on prediction with abstention. Neeman et al. (2023); Zhou et al. (2024) introduced the concept of answerability augmentation, where LLMs are trained to respond with "Unanswerable" when presented with irrelevant or randomly generated contexts. This ensures that the models do not make incorrect predictions in the absence of reliable evidence. Further studies Wang et al. (2023, 2022) have developed confidence calibration techniques, which encourage LLMs to avoid overly confident predictions in ambiguous or uncertain situations.
In our work, we investigate the context-faithfulness of LLMs when faced with conflicting knowledge. We define a model as context-faithful if it demonstrates strong receptiveness to new facts and evidence, adapting its responses accordingly. This capability is essential for ensuring the reliability of LLMs in the RAG system.
@@ 2.2 Construction of Knowledge Conflicts
In controlled experiments, knowledge conflicts are typically simulated by constructing counterfactual memories based on a model’s parametric memory. Various heuristic approaches have been proposed for this purpose, such as negation injection Kassner et al. (2021); Petroni et al. (2020); Pan et al. (2021), which alters facts by introducing negations, and entity substitution Longpre et al. (2021); Chen et al. (2022); Si et al. (2023); Zhou et al. (2023), which replaces mentions or entities in the parametric memory with alternatives to generate counter-memory answers. However, these techniques are constrained to word-level edits, which can lead to low coherence across the constructed counter-memory. To address this limitation, recent studies Xie et al. (2024); Jin et al. (2024) have explored generating counter-memories using LLMs, producing more coherent and consistent counterfactual content. We adopt this approach in generating our dataset, ensuring the counter-memories maintain a higher level of coherence.
@@ 3 Methodology
Figure 1: Framework for Evaluating LLMs’ Evidence Context-faithfulness. In Step 1, we calculate the memory strength of a question by calculating the consistency of answers to different paraphrases of the same question. In Step 2, we generate MA (memory answer) with closed-book setting and CMA (counter memory answer) by modifying answer entity in MA. Key information should not changed. Step 3 involves generating supporting evidence for the CMA, either as direct or indirect evidence. In Step 4 (not shown), we test the LLM’s response by presenting it with the evidence to evaluate whether it adheres to the CMA or defaults to its MA. All experiments are implemented under the zero-shot setting to avoid the bias introduced by demonstrations.
@@ 3.2 Datasets
  • popQA Mallen et al. (2023) is an entity-centric question-answering dataset comprising 14,000 questions. The dataset is derived from knowledge triples in Wikidata, where questions are generated using question templates specific to different relationship types. popQA aims to capture a realistic, long-tail distribution of entity popularity, making it a valuable resource for studying the performance of lesser-known entities. Xie et al. (2024) use popQA to test the receptiveness of LLMs by eliciting high-quality parametric memory from LLMs and constructing the corresponding counter-memory. We reuse MA and CMA generated by Xie et al. (2024) for our experiments.
  • Natural Questions Kwiatkowski et al. (2019) is widely used in open-domain QA research. It consists of manually crafted questions based on selected paragraphs from Wikipedia, and the subjects in questions of the NQ dataset are generally more popular and commonly known. Longpre et al. (2021) provide a test set that is used to test the context-faithfulness of LLMs by substituting entity of the NQ dataset. The entity substitute involves five categories: person (PER), date (DAT), numeric (NUM), organization (ORG), and location (LOC). The test set contains 4,685 samples, including 1,667 unique questions.
@@ 3.3 Memory Strength
Inspired by Zhao et al. (2024), we use the consistency of answers to different paraphrases of the same question $Q$ to measure the LLM’s memory strength $S_{Q}$ for the knowledge $K_{Q}$ associated with the question. This method is motivated by the intuition that if an LLM does not have a strong memory of a question, it often produces varying answers when presented with different paraphrases but semantically equivalent questions, as shown in Table 8 in Appendix. In contrast, it can produce consistent answers if the LLM has a strong memory of a question. The process involves two key steps: First, several paraphrased versions of the original question are generated with ChatGPT [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo, the specific version is 0125., and the answers to those paraphrased questions are clustered (Section 3.3.1). Then, memory strength $S_{Q}$ is calculated using answer consistency (Section 3.3.2).
@@ 3.3.1 Question Paraphrases and Answer Clustering
The prompt used for paraphrasing the question is provided in Tabel 9 (index 1) in the Appendix. For each question $Q$, we generate $n$ paraphrases $\{P_{1},\cdots,P_{n}\}_{Q}$. For the NQ dataset, we paraphrase the question in each data sample directly. For the popQA dataset, we paraphrase the question template for each relation type since all questions of the same relation type share the same question template. To ensure the paraphrased questions are proper to use, we check if two paraphrased questions are semantically equivalent with an LLM [^2] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B. The prompt for this semantic equivalence detection is provided in Table 9 (index 2). For any paraphrase that is deemed unequivalent, we ask the LLM to re-generate it until a satisfactory version is produced.
@@ 3.3.2 Calculating Memory Strength
Once answer clusters $\{c_{1},\cdots,c_{m}\}_{Q}$ are identified, memory strength $S(Q)$ can be obtained by calculating the entropy of cluster distribution. The formula is
@@
where $N(c_{i})$ is the number of answers in the cluster $c_{i}$, and $n$ is the number of paraphrases for question $Q$. In the experiments, we set $n=7$ for all the questions in the NQ and popQA datasets. Memory strength reflects how well the LLM remembers the required knowledge: the weaker the memory, the more random and inconsistent the answers are.
@@ 3.4.1 MA and CMA Generation
For the popQA dataset, both MA and CMA are obtained following the method described in Xie et al. (2024). For the MA of the NQ dataset, we also use a closed-book approach, similar to Xie et al. (2024). While, the process for generating CMA differs. Unlike the popQA dataset, the NQ dataset does not provide relation types for the questions or offer sets of subject and object entities for substitution. To address this issue, we propose an approach using an LLM to substitute entities in MA to generate CMA. First, we identify which “wh-” question type333which refers to what, when, where, who, whom, which, whose, why, and how. the question belongs to using string matching. Then, based on the question type, we determine the type of entity to be replaced in the MA. Finally, we use an LLM to make the substitution. For example, in Figure 1, the question “how many episodes…” is of the type “how_many”, so the entity to be replaced in the MA “there are 23 episodes…” should be a NUMBER. We let ChatGPT perform the substitution with an alternative entity. The prompt used is shown in Table 9 (index 5). We have the detailed description for generating CMA in Appendex B.1.
CMA filter. As noted in Section 3.3, LLMs can produce multiple MAs for the same question. To ensure the CMA conflicts with MAs, we require that the CMA is different from any of the answers $\{A_{1},\cdots,A_{n}\}$ generated in Section 3.3.1, so the alternative entity should not appear in MAs. For the popQA dataset, the alternative entity is known. For NQ dataset, we first identify the alternative entity in the CMA by comparing the MA and CMA, and then check if this entity appears in any of the MAs $\{A_{1},\cdots,A_{n}\}$. We filter out data samples whose CMA does not conflict with MAs.

### v2

# Investigating Context Faithfulness in Large Language Models: The Roles of Memory Strength and Evidence Style

@@ Abstract
Retrieval-augmented generation (RAG) improves Large Language Models (LLMs) by incorporating external information into the response generation process. However, how context-faithful LLMs are and what factors influence LLMs’ context faithfulness remain largely unexplored. In this study, we investigate the impact of memory strength and evidence presentation on LLMs’ receptiveness to external evidence. We quantify the memory strength of LLMs by measuring the divergence in LLMs’ responses to different paraphrases of the same question, which is not considered by previous works. We also generate evidence in various styles to examine LLMs’ behavior. Our results show that for questions with high memory strength, LLMs are more likely to rely on internal memory. Furthermore, presenting paraphrased evidence significantly increases LLMs’ receptiveness compared to simple repetition or adding details. These findings provide key insights for improving retrieval-augmented generation and context-aware LLMs. Our code is available at https://github.com/liyp0095/ContextFaithful.
Investigating Context Faithfulness in Large Language Models: The Roles of Memory Strength and Evidence Style
Yuepei Li,   Kang Zhou,   Qiao Qiao,   Bach Nguyen,   Qing Wang,   Qi Li Department of Computer Science Iowa State University, Ames, Iowa, USA {liyp0095, kangzhou, qqiao1, ntbach, qingwang, qli}@iastate.edu
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Fan et al. (2024); Zhao et al. (2023) has gained increasing popularity as it improves the performance of Large Language Models (LLMs) by integrating external information during the generation process, particularly when the model’s internal knowledge is insufficient or outdated Bianchini et al. (2024); Procko (2024); Siriwardhana et al. (2023); Vakayil et al. (2024); Wang et al. (2024); Jeong (2023). It raises the importance of the study of how context-faithful LLMs are. In this study, we explore whether LLMs are context-faithful when encountering external information, particularly when that information conflicts with the LLMs’ internal memory.
Figure 1: Demonstration of issues in context faithfulness testing schema: a) LLMs may fail to comprehend long contexts, resulting in low receptiveness to the context. b) Evaluating LLMs on the same dataset may be unfair due to variations in their knowledge.
@@
To investigate the issue of context faithfulness, there are two main approaches to creating knowledge conflict contexts. One approach Longpre et al. (2021); Chen et al. (2022) is entity substitution, which replaces the gold entity in context with a similar one. Another approach Xie et al. (2024); Jin et al. (2024) involves generating counter-memory evidence with LLMs, and these studies have shown that LLMs are generally receptive to external evidence as long as it is coherent.
These methods, however, overlook some important aspects of the task. First, previous work Longpre et al. (2021); Xie et al. (2024) provides long contexts to LLMs, which can be challenging for LLMs to understand Xie et al. (2024). It makes the test difficult to distinguish whether the cause of LLM behavior is due to knowledge conflicts or lacking comprehension ability. For example, in Figure 1 (a), the LLM may overlook both contexts, but the reason why Context 1 is ignored could be a lack of comprehension rather than knowledge conflicts. Second, different LLMs are trained with different data and are likely to obtain different knowledge. Thus, testing LLMs on the same dataset may be unfair. The test may favor LLMs with less knowledge. As shown in Figure 1 (b), LLMs with strong memory are less likely to be correct.
@@
To address these issues, we introduce a method to quantify the memory strength of LLMs and generate short evidence in various styles to examine LLMs’ behavior. Inspired by Zhao et al. (2024), we assess memory strength by measuring the divergence in LLMs’ responses to different paraphrases of the same question. Intuitively, an LLM demonstrates high memory strength when it consistently provides the same answer across all paraphrased versions of a question. For evidence styles, we classify them into direct and indirect forms: direct evidence provides a straightforward answer to the question, while indirect evidence incorporates additional details to support the answer. Unlike Wan et al. (2024) that focuses on how different types of evidence influence LLMs’ behavior, we study how the presentation of evidence affects its convincing availability. Through these methods, we analyze the relationship between context faithfulness and LLM memory strength, and we explore the impact of different evidence styles on context faithfulness. Our conclusions are as follows:
@@
  • The receptiveness of LLMs to external evidence is strongly correlated with memory strength to the question. We observed this relationship both across different datasets and different LLMs. Contrary to the findings of Xie et al. (2024); Jin et al. (2024) that LLMs are highly receptive (less than 5%) to external evidence when it is coherent, we find that the probability of the model relying on its internal memory is non-negligible for questions that the LLMs have a strong memory. For example, GPT-4, which has strong memory on the NQ dataset, answers almost 50% of the questions with internal memory. We also demonstrate an urgent need for memory strength-aware evaluation metrics.
@@
  • The style of the evidence plays an important role in LLM’s receptiveness to external information. Our research demonstrates that presenting the LLM with multiple paraphrases of the same evidence substantially increases its receptiveness. This approach outperforms simple repetition of the evidence and is more effective than adding additional details to the evidence. These findings provide valuable insights to the research of RAG.
@@ 2 Related Work
Origin Question Paraphrased Questions Answers with LLaMA2-7B
Table 1: LLMs give different answers when asked with paraphrased questions.
@@ 2.1 Context Faithfulness of LLM
To update static factual knowledge Lazaridou et al. (2021); Karpukhin et al. (2020); Kasai et al. (2023) in LLMs, the retrieval-based method has been introduced to involve external information to LLMs Lazaridou et al. (2022); Izacard et al. (2024); Khattab et al. (2022); Santhanam et al. (2022); Gao and Callan (2022). However, these methods can introduce knowledge conflicts between the introduced external information (context) and pre-existing internal memories from LLMs. LLMs often rely on their internal memories, and overlook the contextual evidence Longpre et al. (2021). To make LLMs more faithful to context, recent studies Neeman et al. (2023); Li et al. (2023) fine-tune LLMs on counterfactual contexts, where the original facts are replaced with counterfactual ones. Another work Zhou et al. (2023) proposes a novel approach using prompting to improve context faithfulness in LLMs without additional fine-tuning. Recent works Farahani and Johansson (2024); Wadhwa et al. (2024) found that models like T5, LLaMA2-7B, and Phi2-2.7B are generally context-faithful. However, they did not evaluate newer models (e.g., LLaMA3.2, GPT-4) or consider the impact of memory strength, limiting the generality of their conclusions.
A related area of research focuses on prediction with abstention. Neeman et al. (2023); Zhou et al. (2024) introduces answerability augmentation, where LLMs are trained to respond with "Unanswerable" when presented with irrelevant or randomly generated contexts. This ensures that the models do not make incorrect predictions without reliable evidence. Further studies Wang et al. (2023, 2022) develop confidence calibration techniques to improve context faithfulness by encouraging LLMs to avoid overly confident predictions in ambiguous or uncertain situations.
In our work, we investigate the context faithfulness of LLMs when faced with conflicting knowledge. We define a model as context-faithful if it demonstrates high receptiveness to new facts and evidence with strong conflicting memories. This capability is essential to ensure the reliability of LLMs in the RAG system.
@@ 2.2 Construction of Knowledge Conflicts
In controlled experiments, knowledge conflicts are typically simulated by constructing counterfactual memories based on a model’s internal memory. Various heuristic approaches have been proposed for this purpose, such as negation injection Kassner et al. (2021); Petroni et al. (2020); Pan et al. (2021) and entity substitution Longpre et al. (2021); Chen et al. (2022); Si et al. (2023); Zhou et al. (2023). Negation injection alters facts by introducing negations and entity substitution replaces mentions or entities in the evidence with alternatives to generate counter-fact evidence. However, these techniques are constrained to word-level edits, which can lead to low coherence across the constructed evidence. To address this limitation, recent studies Xie et al. (2024); Jin et al. (2024) have explored generating evidence using LLMs, producing more coherent and consistent counterfactual content. We adopt this approach in generating our dataset, ensuring the generated evidence maintains a higher level of coherence.
@@ 3 Methodology
Figure 2: Framework for Evaluating LLMs’ Context Faithfulness. In Step 1, we calculate the memory strength of a question using the consistency of answers to different paraphrases of the same question. In Step 2, we generate MA (memory answer) under the closed-book setting and CMA (counter memory answer) by modifying answer entity in MA while keeping the other information. In Step 3, we generate supporting evidence for the CMA in various styles. In Step 4 (not shown), we test the LLM’s response by presenting the questions with the evidence. All experiments are implemented under the zero-shot setting to avoid the bias introduced by demonstrations.
@@ 3.2 Datasets
  • popQA Mallen et al. (2023) is an entity-centric question-answering dataset comprising 14,000 questions. The dataset is derived from knowledge triples in Wikidata, where questions are generated using question templates specific to different relationship types. popQA aims to capture a realistic, long-tail distribution of entity popularity, making it a valuable resource for studying the performance of lesser-known entities. Xie et al. (2024) use popQA to test the receptiveness of LLMs by eliciting high-quality internal memory from LLMs and constructing the corresponding counter-memory. We reuse MA and CMA generated by Xie et al. (2024) for our experiments.
  • Natural Questions Kwiatkowski et al. (2019) is widely used in open-domain QA research. It consists of manually crafted questions based on selected paragraphs from Wikipedia, and the subjects in questions of the NQ dataset are generally more popular and commonly known. Longpre et al. (2021) provide a test set that is used to test the context faithfulness of LLMs by substituting entity of the NQ dataset. The entity substitute involves five categories: person (PER), date (DAT), numeric (NUM), organization (ORG), and location (LOC). The test set contains 4,685 samples, including 1,667 unique questions.
@@ 3.3 Memory Strength
Inspired by Zhao et al. (2024), we use the consistency of answers to different paraphrases of the same question $Q$ to measure the LLM’s memory strength $S_{Q}$ for the knowledge $K_{Q}$ associated with the question. This method is motivated by the intuition that if an LLM does not have a strong memory of a question, it will often give different answers when asked with paraphrased questions that are semantically equivalent, as shown in Table 1. In contrast, it can produce consistent answers if the LLM has a strong memory of a question. The process involves two key steps: First, several paraphrased versions of the original question are generated with ChatGPT [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo, the specific version is 0125., and the answers to those paraphrased questions are clustered (Section 3.3.1). Then, memory strength $S_{Q}$ is calculated using answer consistency (Section 3.3.2).
@@ 3.3.1 Question Paraphrases and Answer Clustering
The prompt used for paraphrasing the question is provided in Tabel 11 (index 1) in the Appendix. For each question $Q$, we generate $n$ paraphrases $\{P_{1},\cdots,P_{n}\}_{Q}$. For the NQ dataset, we paraphrase the question in each data sample directly. For the popQA dataset, we paraphrase the question template for each relation type since all questions of the same relation type share the same question template. To ensure the paraphrased questions are proper to use, we check if two paraphrased questions are semantically equivalent with an LLM [^2] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B. The prompt for this semantic equivalence detection is provided in Table 11 (index 2). For any paraphrase that is deemed not equivalent, we ask the LLM to re-generate it until a satisfactory version is produced.
@@ 3.3.2 Calculating Memory Strength
Once answer clusters $\{c_{1},\cdots,c_{m}\}_{Q}$ are identified, memory strength $S(Q)$ can be obtained by calculating the negative entropy of cluster distribution. The formula is
@@
where $N(c_{i})$ is the number of answers in the cluster $c_{i}$, and $n$ is the number of paraphrases for question $Q$. The memory strength score is a non-positive value. A larger score indicates a stronger memory (0 is the maximum value of memory strength score). In the experiments, we set $n=7$ for all the questions in the NQ and popQA datasets. Memory strength reflects how well the LLM remembers the required knowledge: the weaker the memory, the more random and inconsistent the answers are.
@@ 3.4.1 MA and CMA Generation
For the popQA dataset, both MA and CMA are obtained following the method described in Xie et al. (2024). For the MA of the NQ dataset, we also use a closed-book approach, similar to Xie et al. (2024). While, the process of generating CMA differs from the process of generating CMA in Xie et al. (2024). Unlike the popQA dataset, the NQ dataset does not provide relation types for the questions or offer sets of subject and object entities for substitution. To address this issue, we propose an approach using an LLM to substitute entities in MA to generate CMA. First, we identify which “wh-” question type333which refers to what, when, where, who, whom, which, whose, why, and how. the question belongs to using string matching. Then, based on the question type, we determine the type of entity to be replaced in the MA. Finally, we use an LLM to make the substitution. For example, in Figure 2, the question “how many episodes…” is of the type “how_many”, so the entity to be replaced in the MA “there are 23 episodes…” should be a NUMBER. We let ChatGPT perform the substitution with an alternative entity. The prompt used is shown in Table 11 (index 5). We have the detailed description for generating CMA in Appendix A.1.
CMA filter. As noted in Section 3.3, LLMs can produce multiple MAs for the same question. To ensure the CMA conflicts with MAs, we require that the CMA is different from any of the answers $\{A_{1},\cdots,A_{n}\}$ generated in Section 3.3.1, so the alternative entity should not appear in MAs. For the popQA dataset, the alternative entity is known. For the NQ dataset, we first identify the alternative entity in the CMA by comparing the MA and CMA, and then check if this entity appears in any of the MAs $\{A_{1},\cdots,A_{n}\}$. We filter out data samples whose CMA does not conflict with MAs.